use three multipliers to write the expression needed to convert 45 cubic kilometers to cubic meters.

Answer by mananth(16946)

1 km = 1000 m
1cu.km = 1000*1000*1000 cu.m
=10^9 cu.m
45 cu.km = 45*10^9 cu.m
= 4.5 * 10^10 cu.m
...
m.ananth@hotmail.ca